OpenStudy (anonymous):

Help finding all solutions of this equation in the interval [0, 2pi)? (Will give medal & fan) tan^x - sec x = -1

hero (hero):

Are you sure your equation is correct as written?

hero (hero):

There seems to be something missing from the tangent

OpenStudy (anonymous):

tan^2 x - sec x = -1, sorry

hero (hero):

Okay, so basically, what you have to do here is convert the trigonometric equation to a quadratic equation.

OpenStudy (anonymous):

How would I do that?

hero (hero):

Well, if you remember your Pythagorean Identities, you would know that \(\tan^2x = \sec^2x - 1\) right?

OpenStudy (anonymous):

Yeah

hero (hero):

So since that is true, this means \(\tan^2 x - \sec x = -1\) can be re-written as \(\sec^2x - 1 - sec x = -1\)

hero (hero):

Furthermore, If we add 1 to both sides, we'll have \(\sec^2x - \sec x = 0\) Do you agree?

OpenStudy (anonymous):

Yes I agree

hero (hero):

And now we have the trigonometric equation in quadratic form. Do you see what we might be able to do for the next step?

OpenStudy (anonymous):

hmm... factor...?

hero (hero):

Yes, correct. So how would the equation look after factoring?

OpenStudy (anonymous):

(sin x)(sin x) - sin x?

hero (hero):

Hmmm. I thought we were working with sec(x)

OpenStudy (anonymous):

My bad, the same thing but with sec. I'm exhausted lol

hero (hero):

Well, actually, maybe we should let y = sec(x) for now.

hero (hero):

If that's the case, then \(\sec^2(x) - \sec(x) = 0\) would become \(y^2 - y = 0\)

hero (hero):

So factor \(y^2 - y = 0\)

OpenStudy (anonymous):

I'm not really that great at factoring /: (y - 1)(y + 1)?

OpenStudy (anonymous):

Sorry my internet went down for a little. I'm still confused

hero (hero):

\(y^2 - y = 0\) factors to \(y(y - 1) = 0\) Likewise \(\sec(x)(\sec(x) - 1) = 0\) Furthermore \(\sec(x) = 0\) or \(\sec(x) = 1\)